mykenzie:

Martina chose a random sample of 10 bags of candy from two different brands, A and B. All bags in the sample are the same size. Martina counts the number of candy pieces in each bag and plots the results on two dot plots. A dot plot is shown. The title of the plot is Brand A. The line is numbered with tick marks at 40, 45, 50, 55, 60, 65, and 70 and is labeled Number of Candy Pieces. There are 2 dots at 52, 1 dot at 53, 2 dots at 55, 1 dot at 56, 2 dots at 57, and 2 dots at 62. A dot plot is shown. The title of the plot is Brand B. The line is numbered with tick marks at 40, 45, 50, 55, 60, 65, and 70 and is labeled Number of Candy Pieces. There is 1 dot at 44, 1 dot at 46, 2 dots at 52, 1 dot at 55, 1 dot at 62, 1 dot at 65, 2 dots at 67 and 1 dot at 69 Based on the plots, which statement is a valid comparison of the number of candies in the bags of the two brands? A. The number of candies in the bags from brand B is greater and more consistent than the number of candies in the bags from brand A. B. The number of candies in the bags from brand B is greater and less consistent than the number of candies in the bags from brand A. C. The number of candies in the bags from brand B is fewer and more consistent than the number of candies in the bags from brand A. D. The number of candies in the bags from brand B is fewer and less consistent than the number of candies in the bags from brand A.

mykenzie:

1 attachment
mykenzie:

1 attachment
mhanifa:

Compare the ranges on the dot plots, it will give you and idea on consistency

mhanifa:

Compare maximum values, it will answer to the first part of the question

mykenzie:

63

Florisalreadytaken:

2 dots at 52, 1 dot at 53, 2 dots at 55, 1 dot at 56, 2 dots at 57, 2 dots at 62.


10 total dots so we can write: \[ \frac{\left(2\times 52+53+2\times 55+56+\ 2\times 57+2\times 62\right)}{10} \] \[ \frac{\left(104+53+110+56+\ 114+124\right)}{10} \] \[ \frac{561}{10} \] \[ \approx 56.1 \] 1 dot at 44 1 dot at 46 2 dots at 52 1 dot at 55 1 dot at 62 1 dot at 65 2 dots at 67 1 dot at 69
10 total dots as well: \[ \frac{44+46+2\times 52+55+62+65+2\times 67+69}{10} \] \[ \frac{44+46+104+55+62+65+134+69}{10} \] \[ \frac{579}{10} \] \[ \approx57.9 \] So by comparing the two, we can see that company b provides more candies in their bags all thats left is to have a look at what the graphs look like, as you know that the number of candies in bags B is greater than the bags A. can you pull the answer?